### v1

# Requirements Coverage-Guided Minimization for Natural Language Test Cases

Abstract.
As software systems evolve, test suites tend to grow in size and often contain redundant test cases. Such redundancy increases testing effort, time, and cost. Test suite minimization (TSM) aims to eliminate such redundancy while preserving key properties such as requirement coverage and fault detection capability. In this paper, we propose RTM (Requirement coverage-guided Test suite Minimization), a novel TSM approach designed for requirement-based testing (validation), which can effectively reduce test suite redundancy while ensuring full requirement coverage and a high fault detection rate (FDR) under a fixed minimization budget. Based on common practice in critical systems where functional safety is important, we assume test cases are specified in natural language and traced to requirements before being implemented. RTM preprocesses test cases using three different preprocessing methods, and then converts them into vector representations using seven text embedding techniques. Similarity values between vectors are computed utilizing three distance functions. A Genetic Algorithm, whose population is initialized by coverage-preserving initialization strategies, is then employed to identify an optimized subset containing diverse test cases matching the set budget.
We evaluate RTM on an industrial automotive system dataset comprising $736$ system test cases and $54$ requirements. Experimental results show that RTM consistently outperforms baseline techniques in terms of FDR across different minimization budgets while maintaining full requirement coverage. Furthermore, we investigate the impact of test suite redundancy levels on the effectiveness of TSM, providing new insights into optimizing requirement-based test suites under practical constraints.
Test Suite Minimization, Requirements Testing, Genetic Algorithm, Automotive System
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: xxx; 2025; XXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Requirements analysis
1. Introduction
Software testing is a critical activity throughout the software development life cycle, playing a key role in early defect detection and compliance with requirements (Myers et al., 2011; Everett and McLeod Jr, 2007). In particular, testing driven by requirements is conducted to validate systems before their deployment (Dustin et al., 2009). This is especially critical for systems that must demonstrate functional safety, such as automotive systems. In such contexts, it is also common to specify test cases in natural language (following templates) to facilitate their analysis and traceability to requirements, before manually writing the corresponding test code.
However, due to the increasing complexity of systems and their continuous change by large teams, test suites are hard to maintain, tend to grow, and are often redundant (i.e., some of the test cases are likely to detect the same faults), thus leading to waste of time and resources (Rothermel et al., 2002). Further, in practice, testers have limited resources and time to test new system updates. To address this, given a test budget, test suite minimization (TSM) is proposed to prune test cases that are most likely to be redundant, while satisfying coverage criteria and maximizing fault detection (Harrold et al., 1993; von Ronne, 1999), thus significantly reducing testing cost and time while limiting the impact on system quality or compliance.
While a number of approaches have been proposed for TSM (Noemmer and Haas, 2019; Miranda and Bertolino, 2017; Coviello et al., 2018; Liu et al., 2011; Viggiato et al., 2022; Cruciani et al., 2019; Zhang et al., 2019; Hemmati et al., 2013; Pan et al., 2023, 2024; Marchetto et al., 2017; Xia et al., 2021; Anwar et al., 2019; Yoo and Harman, 2010), most existing techniques rely on code and structural coverage criteria (e.g., statement or branch coverage). These techniques do not address the unique challenges of minimizing test suites driven by system requirements, with test cases often specified in natural language. Moreover, to the best of our knowledge, no existing solution is capable of simultaneously maintaining full requirement coverage and adhering to a fixed minimization budget, a common situation in industrial practice due to limited time and resources.
To bridge this gap, we propose RTM (Requirements coverage-guided Test suite Minimization), a novel approach specifically designed for requirement-based testing. RTM aims to reduce test suite redundancy while ensuring full requirement coverage, making it particularly well-suited to meet the practical needs of our industrial partners, and critical systems that must demonstrate functional safety in general. RTM takes requirement-based test cases–composed of natural language test steps–as input and computes pairwise similarity between them. Specifically, we investigate three different preprocessing methods to normalize the test steps, followed by seven distinct text embedding techniques–both sentence-level and word-level–to transform the textual data into vector representations: Term Frequency–Inverse Document Frequency (TF-IDF), Universal Sentence Encoder (USE) (Cer et al., 2018), LongT5 (Guo et al., 2022), Amazon Titan Text Embedding V2 (AWS, 2024), Word2Vec (Mikolov et al., 2013), GloVe (Pennington et al., 2014), FastText (Bojanowski et al., 2017). We then measure the similarity between test cases using three different distance metrics: cosine similarity and Euclidean distance for sentence-level embeddings, and Word Mover’s Distance (WMD) (Kusner et al., 2015) for word-level embeddings. Based on the computed similarities, we employ a Genetic Algorithm (GA) (Luke, 2013) to search for the optimal test suite that minimizes similarity, adheres to a fixed budget, and ensures full requirement coverage. To help the GA efficiently find valid subsets, we devise three initialization strategies designed to generate an initial population consisting of subsets satisfying both coverage and budget constraints.
RTM was evaluated on a real-world automotive dataset consisting of $736$ system test cases across seven test runs, along with $54$ requirements. The results demonstrate that RTM significantly outperforms two baseline techniques in terms of fault detection rate (FDR) under various minimization budgets. Furthermore, we explore the impact of the redundancy level of the test suites on FDR. Our findings show that: ① RTM consistently outperforms all baseline techniques across different redundancy levels; ② The level of redundancy has a strong effect on the achieved FDR, for all TSM techniques; ③ Especially at lower minimization budgets and for highly redundant test suites, there is still significant room for improvement between RTM and the theoretical maximum FDR.
In summary, our contributions include:
  • We propose RTM, a TSM approach tailored for requirement-based testing, where test cases are specified in natural language. RTM ensures full requirement coverage while operating under a fixed minimization budget, addressing a practical constraint often overlooked in prior work.
  • We validate RTM on automotive system requirements and test cases, and compare it with a number of baseline techniques for test suite minimization across varying minimization budgets, demonstrating its superior fault detection capability.
  • We further investigate the impact of redundancy levels of test suites on the effectiveness of TSM, offering new insights into the relationship between test suite redundancy and TSM fault detection performance.
  • To foster reproducibility and facilitate further research, we have open-sourced our replication package—including source code and experimental scripts—which is publicly available at https://anonymous.4open.science/r/RTM-5C98/. Industrial requirements and test cases are confidential.
The rest of the paper is organized as follows: Section 2 defines the TSM problem and outlines the context of this study. Section 3 describes our approach: RTM. The study design is detailed in Section 4, followed by experimental results in Section 5. Section 6 discusses threats to validity. Section 7 reviews related work. Finally, Section 8 concludes the study with a discussion of future work.
2. Problem Definition: Test Suite Minimization in Automotive Requirements Testing
In this paper, we aim to minimize test suites derived from automotive system requirements. In this section, we first describe the format and structure of the test cases in our dataset, followed by a formal definition of the TSM problem in our specific context.
2.1. Test Cases based on Automotive Requirements
Like other critical systems, modern automotive software systems are expected to comply with functional safety standards such as ISO 26262 (Siegl et al., 2010). To ensure compliance, system engineers derive and maintain extensive test suites based on system requirements, which are executed to validate system behavior against expected requirements. Typically, test cases are specified in natural language, following templates, and then manually converted into test code and implemented by experts.
The test cases in our dataset are collected from our industry partner and designed for testing automotive system requirements. As shown in Figure 1, these test cases are specified in natural language. Each test case consists of a sequence of diagnostic actions, structured into clearly labeled steps. These steps typically include setting up the preconditions for testing, creating the fault conditions, checking how the automotive system detects the issues through Diagnostic Trouble Codes (DTCs), and confirming the system properly clears the DTC. The detailed test actions involve reading and setting system variables, awaiting signal conditions, sending diagnostic requests, and checking system responses. To this end, the test cases exhibit the following characteristics:
Structured and Precise: Test cases explicitly define preconditions, test steps, and expected outcomes.
Domain-Specific Vocabulary: Test cases extensively utilize automotive-specific terminology, signals, variables, and abbreviations.
However, as the number of system requirements increases, the corresponding test suites can become excessively large, leading to dramatically increased testing costs. Moreover, as new requirements are added, test cases are often written without checking the existing ones. This can lead to functional overlap between test cases, thus increasing redundancy in the test suite.
STEP 1 Set Global Preconditions Read variable Variable_A STEP 2 Set Valid Preconditions Set System variable Variable_1 = 1 Await Value Match Signal SIGNAL_A = 1 STEP 3 Create Fault Condition Set System variable Variable_2 = 1 STEP 4 Verify DTC Maturation Time Check maturation time (Expected: X ms) STEP 5 Check the DTC is Active Read variable Variable_A Send request PATH_TO_REQUEST_A Check expected diagnostic response Set System variable Variable_1 = 0 Await Value Match Signal Variable_1 = 0 STEP 6 Remove DTC Condition Set System variable Variable_2 = 0 STEP 7 Verify DTC Dematuration Time Read variable Variable_A Check dematuration time (Expected: X ms) Send request PATH_TO_REQUEST_A Check expected diagnostic response
Figure 1. A sanitized example of the test case in our dataset
2.2. Test Suite Minimization
Test suite minimization (TSM) aims to reduce the size of the test suite while ensuring important testing objectives such as requirement coverage, fault detection capability, and test case diversity (Harrold et al., 1993; von Ronne, 1999). Given that test suite diversity is reported to be positively correlated with fault detection capability (Hemmati et al., 2013), minimizing similarity among test cases can enhance the fault detection capability of the minimized test suite. When applied properly, TSM can significantly reduce testing costs and time while minimizing its impact on system quality or compliance.
Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ be the set of system requirements for an automotive system, and let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{m}\}$ be the corresponding test suite. Each test case $t_{i}\in\mathcal{T}$ covers a subset of requirements denoted by $Cover(t_{i})\subseteq\mathcal{R}$. Note that in our dataset, each test case only covers one requirement, whereas each requirement can be covered by many test cases. This is because each test case is specifically designed to validate one scenario where the requirement is expected to hold. Moreover, covering all requirements is a must in automotive testing, to comply with functional safety standards (Siegl et al., 2010), which is also strictly required by our industry partners. Let $Sim(t_{i},t_{j}):\mathcal{T}\times\mathcal{T}\rightarrow[0,1]$ denote the similarity between test cases $t_{i}$ and $t_{j}$.
Given a test budget $\mathcal{M}\in(0,1]$, the goal is to select a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ such that:
  • Coverage: $\bigcup_{t\in\mathcal{T}^{\prime}}Cover(t)=\mathcal{R}$,
  • Budget: $|\mathcal{T}^{\prime}|=|\mathcal{T}|\times\mathcal{M}$,
  • Diversity: the total pairwise similarity among test cases in $\mathcal{T}^{\prime}$ is minimized.
Formally, this can be defined as the following optimization problem:
Given: test suite $\mathcal{T}$, requirement set $\mathcal{R}$, budget $\mathcal{M}\in(0,1]$,
coverage function $Cover(t)\subseteq\mathcal{R}$ for each $t\in\mathcal{T}$, and similarity function $Sim(t_{i},t_{j})$.
Find: a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ such that:
(1) $|\mathcal{T}^{\prime}|$ $=|\mathcal{T}|\times\mathcal{M},$
(2) $\bigcup_{t\in\mathcal{T}^{\prime}}Cover(t)$ $=\mathcal{R},$
(3) $\sum_{\begin{subarray}{c}t_{i},t_{j}\in\mathcal{T}^{\prime}\\ i<j\end{subarray}}Sim(t_{i},t_{j})$ is minimized.
This problem extends the classical set cover problem, which is known to be NP-hard, by introducing two additional constraints: a strict minimization budget and a diversity objective. In particular, the selected test subset must maintain full requirement coverage under a fixed budget while minimizing the overall similarity among test cases. As a result, solving this problem exactly is computationally intractable for large-scale industrial systems. Therefore, practical heuristics or approximation strategies are typically employed in automotive domains to achieve efficient and effective test suite minimization, especially when additional factors, such as fault detection capability, execution cost, or coverage redundancy, are also considered.
3. Approach
This section introduces RTM, a TSM approach that satisfies the three criteria outlined in Section 2.2. It targets test cases derived from system requirements, for the purpose of validation. In short, it leverages natural language processing (NLP) techniques and large language models (LLMs) as well as various distance functions to compute test case similarity and then employs evolutionary search to optimize minimization.
As illustrated in Figure 2, RTM first preprocesses test cases using three different strategies, then converts them into vector representations using a variety of text embedding techniques, including both sentence-level and word-level representations: TF-IDF, USE, LongT5, Amazon Titan Text Embedding V2, Word2Vec, GloVe, and FastText. Different distance functions are employed to compute the similarity values between test case embeddings, which are then used as guidance for the evolutionary search to find the optimal minimized test suite $\mathcal{T}^{\prime}$ with diverse test cases. The primary objective of RTM is to find the optimal $\mathcal{T}^{\prime}$ that maximizes fault detection capability by minimizing the similarity values between the test cases while reducing the testing cost, measured in terms of test suite size, as well as ensuring $100\%$ requirement coverage.
[…]
Figure 2. Overall Framework of RTM.
3.1. Test Case Preprocessing
Given that the test cases are expressed in natural language, different preprocessing strategies may affect the outcomes of the text embedding techniques. To investigate this, we implemented three levels of preprocessing:
[…]
  • Preprocessing Method 2 ($PM2$) extends $PM1$ by further stripping punctuation and performing tokenization and lemmatization.
  • Preprocessing Method 3 ($PM3$) serves as a control setting, where no preprocessing is applied, and the raw test case content is used as-is.
3.2. Test Case Representation
After the preprocessing step, we transform the test cases into vector representations using various embedding techniques. We employ both word-level and sentence-level embedding techniques to capture semantic information from test cases. For word-level embeddings, we use Word2Vec, GloVe, and FastText, each of which converts every word in the test case to a $300-$dimensional vector.
For sentence-level embeddings, we employ TF-IDF, USE, LongT5, and Amazon Titan Embedding V2, each of which converts the entire test case content to a fixed-length vector representation that captures its overall semantic meaning, with the dimensionality varying depending on the embedding technique used.
3.3. Similarity Measure
To assess the similarity between test case pairs, we utilize different distance metrics based on the type of embedding used.
For word-level embeddings (i.e., Word2Vec, GloVe, and FastText), we use WMD (Kusner et al., 2015), which measures the minimum cumulative distance required to align the words in one test case with those in another. WMD leverages the underlying word embedding space and is particularly effective in capturing semantic similarity when the word order or exact word match is less important (Kusner et al., 2015).
For sentence-level embeddings (i.e., TF-IDF, USE, LongT5, and Amazon Titan Embedding V2), we applied both cosine similarity and Euclidean distance, which have been widely used in machine learning, information retrieval, and NLP tasks to measure similarity or dissimilarity between vectors (Gomaa and Fahmy, 2013; Huang et al., 2008; Li and Han, 2013; Salton and Buckley, 1988). However, they operate differently and are suited for different use cases: cosine similarity measures the angle between two vectors while ignoring the magnitude, whereas Euclidean distance measures the straight-line distance between two vectors. By applying different types of similarity metrics, we aim to explore how different vector representations and distance functions affect the ability of RTM to detect redundant test cases.
This step constructs a similarity matrix by computing the pairwise similarity between test cases. It takes the vector representations of test case pairs as input and calculates the similarity scores to quantify their similarity.
3.4. Evolutionary Search for Test Suite Minimization
3.4.1. Problem Formalization
Given the test suite before minimization $\mathcal{T}$, the minimization problem is defined as a fixed-size subset selection with $100\%$ requirement coverage as the constraint. Each solution (i.e., subset $\mathcal{T}^{\prime}$) is represented as a binary vector:
(4) $\mathbf{x}=[x_{1},x_{2},\ldots,x_{m}],\quad x_{i}=\begin{cases}1,&\text{if test case }i\text{ is selected}\\ 0,&\text{otherwise}\end{cases}$
where $m$ denotes the total number of test cases in the test suite before minimization (i.e., $\mathcal{T}$).
Given the minimization budget $\mathcal{M}\in(0,1]$, the number of test cases in the subset $\mathcal{T}^{\prime}$ is constrained by:
(5) $\sum_{i=1}^{m}x_{i}=\mathrm{round}\left(m\cdot\mathcal{M}\right)$
where $\mathrm{round}()$ denotes rounding the number to the nearest integer.
Moreover, the test cases in the subset must cover all the requirements:
(6) $\bigcup_{\{i\mid x_{i}=1\}}Cover(i)=\mathcal{R}$
where $Cover(i)$ is the requirement covered by test case $i$ and $\mathcal{R}$ is the set of requirements covered by the test suite before minimization (i.e., $\mathcal{T}$).
Given that the objective of the search process is to minimize the similarity values between the pairs of selected test cases, the fitness function is formulated as:
(7) $Fitness=\frac{\sum_{i,t_{i}\in\mathcal{T}^{\prime}_{n}}(Max_{i,j,t_{i},t_{j}\in\mathcal{T}^{\prime}_{n},i\neq j}\ Norm\_Sim(t_{i},t_{j}))^{2}}{n}$
where $\mathcal{T}^{\prime}_{n}$ denotes the subset containing $n$ test cases, and $Norm\_Sim(t_{i},t_{j})$ represents the normalized similarity score between test cases pair $(t_{i},t_{j})$.
3.4.2. Initialization strategies
Our minimization problem is defined with two hard constraints: (1) the percentage of selected test cases in the subset should satisfy the minimization budget, and (2) the subset should preserve $100\%$ requirement coverage. In order to efficiently guide the search to find valid subsets, we devise three different strategies to initialize a set of valid subsets for the GA to use as a base for further search. Note that in our dataset, the relationship between test cases and requirements is many-to-one. Consequently, one limitation of the minimization problem with full requirement coverage is that the pre-defined number of selected test cases must not be smaller than the total number of requirements.
  • Strategy 1. Iterative Selection This strategy iteratively selects one test case randomly under each requirement, continuing this process until the total number of selected test cases reaches the minimization budget.
  • Strategy 2. Initial Requirement + Random
    • Step 1: Randomly select one test case for each requirement.
    • Step 2: Randomly select the remaining test cases without considering requirements until the total number of selected test cases reaches the minimization budget.
  • Strategy 3. Proportional Selection
    • Step 1: Randomly select a proportion of test cases (approximately) equal to the minimization budget under each requirement.
    • Step 2: If the total number of selected test cases across all requirements is less than or greater than the minimization budget, iteratively add or reduce the number of selected test cases per requirement, ensuring every requirement is still covered, until the desired minimization budget is reached.
3.4.3. Evolutionary Search
After initializing a population of valid subsets using the strategies described in Section 3.4.2, we apply a GA to iteratively evolve the population toward solutions that minimize redundancy while preserving full requirement coverage and complying with the minimization budget.
Each individual in the population is encoded as a binary vector, as defined in Equation 4. The fitness of an individual is computed using the objective function described in Section 3.4.1.
At each generation, we perform the following operations:
  • Selection: We used a binary tournament selection operator that prioritizes valid subsets—those that cover all requirements—over invalid ones. When both candidate subsets are valid, the one with the lower fitness value is selected.
  • Crossover: We used the customized crossover operator (Blank and Deb, 2020) that ensures the offspring maintains a fixed subset size. It first includes all test cases shared by both parent subsets, then adds the test cases present in either parent subset until the minimization budget is met.
  • Mutation: We used an inversion mutation operator (Blank and Deb, 2020) that randomly selects a segment of the subset and reverses its order. Since each subset is represented as a binary vector, this approach preserves the subset size.
Note that crossover and mutation may produce invalid individuals (i.e., those that violate the coverage constraint). To address this, we experimented with incorporating a repair operator during the search process. The repair operator randomly adds one test case for each uncovered requirement and removes the test cases under requirements that are covered by more than one test case, therefore reestablishing full coverage while maintaining the budget constraint for each solution. However, the results obtained using the repair operator were slightly worse than those without it. This can be attributed to the fact that the repair operator reduces the diversity of the solutions. Therefore, we decided not to use the repair operator in the subsequent experiments.
We follow the published guidelines  [^1] https://www.obitko.com/tutorials/genetic-algorithms/recommendations.php used in prior work (Pan et al., 2024) for setting GA hyperparameters. Specifically, we use a population size of $100$, a mutation rate of $0.01$, and a crossover rate of $0.90$. For each generation, the size of each solution (i.e., subset) remains fixed and equal to the minimization budget. The evolutionary process is repeated until the convergence, which is defined as the improvement in the fitness score being less than $0.0025$ across generations. The final output is the best subset identified during the search, which satisfies all constraints and minimizes the redundancy among selected test cases, indicated by the lowest fitness value.
4. Study Design
4.1. Research Questions
  1. RQ1:
    How does RTM perform regarding TSM under different configurations?
    This RQ aims to identify the optimal configuration of RTM that achieves the best performance (i.e., FDR). Specifically, a RTM configuration entails the selection of a preprocessing method (Section 3.1), an embedding technique (Section 3.2), a similarity measure (Section 3.3), and an initialization strategy (Section 3.4.2).
  2. RQ2:
    How does RTM compare to state-of-the-art TSM techniques?
    This RQ aims to compare the performance of RTM with two baseline approaches, namely Random Minimization and FAST-R (Cruciani et al., 2019), under various minimization budgets (i.e., 10%, 20%, …, 90%).
  3. RQ3:
    How does the redundancy level of the test suite impact the effectiveness of TSM approaches?
    The performance of the TSM technique is influenced not only by the selected configuration but also by the characteristics of the test suites. This RQ aims to investigate and quantify the impact of the test suite redundancy level on the effectiveness (i.e., FDR) of TSM.
4.2. Dataset
We evaluated RTM and compared it to baselines on a dataset collected from our industry partner, an automotive system company. There are $736$ test cases across seven test runs. As shown in Figure 1, the test cases consist of test steps that validate the behavior of the automotive system related to a specific DTC. Each test case covers one requirement, with a total of $54$ requirements covered, and may detect multiple faults, with a total of $220$ unique faults detected.
Note that we identified the faults, which are the root causes of the failing test cases (e.g., the parameter values mismatched, the connection error of specific components), using the test execution logs under the guidance of the system engineers. For each test case, we first extracted the failure messages from the test execution logs and then grouped the ones that were triggered by the same root cause into faults.
Test suites for different redundancy levels. In this dataset, one fault can be detected by multiple test cases, indicating a degree of redundancy in the test suite. We define the redundancy level for the test suite as follows:
(8) $RL=\frac{\sum_{i=1}^{m}t_{f_{i}}}{F_{unique}}$
where $m$ is the total number of test cases in this test suite, $t_{f_{i}}$ is the number of faults detected by the test case $t_{i}$, and $F_{unique}$ is the number of unique faults detected by this test suite.
For example, for a test suite $\mathcal{T}$ with three test cases, each of which detects $3$, $5$, and $3$ faults, respectively. The total number of unique faults detected by $\mathcal{T}$ is $4$, and the redundancy level of $\mathcal{T}$ is therefore $(3+5+3)/4=2.75$. Similarly, the redundancy level of our dataset is $2610/220=11.86$.
In order to assess the impact of the test suite redundancy level on the performance of RTM, we generate $10$ diverse test suites for each redundancy level using Integer Linear Programming (ILP)  [^2] https://pypi.org/project/PuLP/ and GA. This process starts by utilizing ILP to search for $100$ test suites with different test suite sizes while satisfying three constraints: (1) $100\%$ requirement coverage, (2) $100\%$ fault coverage, and (3) a specific test redundancy level. We set $15$ different redundancy levels, ranging from $4.5$ to $11.5$, with an interval of $0.5$. Note that there is no test suite that satisfies a redundancy level below $4.5$ under constraints (1) and (2), and the highest redundancy level is $11.86$, which is observed in the original test suite. Then, to enhance the diversity of the dataset for experiments, for the $100$ test suites under each redundancy level, we employ GA to search for the most diverse $10$ test suites that minimize the sum of the set similarity between test suites, which is defined as the Jaccard Similarity between two test suites:
(9) $Jaccard~{}Similarity=\frac{|T_{i}\cap T_{j}|}{|T_{i}\cup T_{j}|}$
where $|T_{i}$cap $T_{j}|$ denotes the number of test cases in the intersection of test suite $T_{i}$ and $T_{j}$. $|T_{i}\cup T_{j}|$ denotes the number of test cases in the union of test suites $T_{i}$ and $T_{j}$.
4.3. Text Representations
Inspired by prior work on vector representations of requirements (Zhao and Alhoshan, 2025; Abbas et al., 2025) and test cases (Viggiato et al., 2022), we employ a variety of text embedding techniques to convert test cases into vector representations. Specifically, we use TF-IDF, USE, Amazon Titan Text Embedding V2, and LongT5 to generate sentence-level embeddings, while Word2Vec, GloVe, and FastText are used to obtain word-level embeddings. Due to limitations in computational resources, data privacy constraints, and input token length restrictions, we were unable to experiment with models such as BERT (Devlin et al., 2019) and Sentence-BERT (Reimers and Gurevych, 2019).
TF-IDF. TF-IDF has demonstrated promising performance in identifying similar test cases expressed in natural language (Viggiato et al., 2022). We also employed TF-IDF to extract numerical vector representations of the test cases. For each word, we calculated its importance within a single test case relative to its occurrence across all other test cases.
USE (Universal Sentence Encoder) (Cer et al., 2018). USE encodes text into high-dimensional vectors, capturing semantic information at the sentence level, making it particularly effective for tasks such as semantic similarity, text classification, clustering, and information retrieval (Cer et al., 2018).
LongT5 (Guo et al., 2022) LongT5 is an extension of the T5 architecture designed to efficiently handle long sequences (Guo et al., 2022). It incorporates sparse attention and reversible layers to reduce memory usage, making it well-suited for tasks like long-document summarization and retrieval-augmented generation. We selected the local attention mechanism for LongT5 as it yielded a higher FDR for TSM compared to the global attention mechanism.
Amazon Titan Text Embedding V2 (AWS, 2024). For LLM-based embedding, we use Amazon Titan Text Embedding V2, which is provided by our industrial partners. The model takes test case steps as input and produces a $1,024$-dimensional vector representation for test cases.
Word2Vec (Mikolov et al., 2013). Word2Vec has been shown to be effective for identifying similar test cases (Viggiato et al., 2022). In this study, we used the Continuous Bag-of-Words (CBOW) architecture with a window size of $10$ as the configuration for Word2Vec, as this setup yielded the highest FDR result for TSM. This was determined with experiments in which we evaluated both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We trained Word2Vec models on the entire test case corpus and transformed each word into $300$-dimensional numerical vectors.
GloVe (Pennington et al., 2014). GloVe is an unsupervised learning algorithm that leverages global word-word co-occurrence statistics from a corpus to learn word embeddings (Pennington et al., 2014). Its core idea is that the ratio of co-occurrence probabilities can capture meaningful linear substructures within the word vector space. We selected a window size of $2$ for GloVe according to its performance in the hyperparameter tuning experiments conducted with window sizes ranging from $2$ to $10$ in increments of $2$. Same as Word2Vec, we used GloVe to generate $300$-dimensional numerical vector representation for each word in the test cases.
FastText (Bojanowski et al., 2017). FastText is a word embedding model proposed by Bojanowski et al. (Bojanowski et al., 2017), which extends Word2Vec by incorporating subword information. Each word is represented as a combination of character n-grams, enabling the model to better handle rare and out-of-vocabulary words, and to generalize more effectively—particularly in morphologically rich or noisy text. We used the CBOW architecture with a window size of $8$ as the configuration for FastText, according to the results of experiments in which we evaluated both CBOW and Skip-Gram architectures with window sizes ranging from $2$ to $10$ in increments of $2$. We used FastText to convert each word in the test cases into a $300$-dimensional numerical vector.
4.4. Baselines
We compare RTM to Random Minimization and FAST-R (Cruciani et al., 2019).
4.4.1. Random Minimization
Since the goal of TSM is to reduce the number of test cases and, consequently, lower testing costs, the most straightforward approach is to randomly remove test cases to meet the minimization budget (Ali et al., 2010). We compare the performance of RTM against random minimization, with and without considering the requirement coverage constraint. For the version with the constraint, we applied the same selection strategy as GA initialization described in Section 3.4.1. For the version without the constraint, we randomly select $k$ test cases from the test suite while ensuring $k$ adheres to the minimization budget.
4.4.2. FAST-R (Cruciani et al., 2019)
FAST-R is a family of similarity-based TSM approaches, including FAST++, FAST-CS, FAST-pw, and FAST-all. FAST++ and FAST-CS utilize k-means++ clustering (Arthur and Vassilvitskii, 2007) and constructed coresets (Bachem et al., 2018). In contrast, FAST-pw and FAST-all leverage minhashing and locality-sensitive hashing (Leskovec et al., 2014). FAST-R can be applied in two scenarios, namely the fixed-budget and adequate scenario. The fixed-budget version of FAST-R minimizes the test suite using a pre-defined minimization budget, while the adequate version prioritizes preserving the requirement coverage of the minimized test suite over adhering to the budget. One limitation of FAST-R is that it cannot simultaneously guarantee both minimization budget and full requirement coverage. We compare RTM against FAST-R under both scenarios where FAST-R is applicable (i.e., fixed-budget and adequate test suite minimization). For FAST-R, we relied on the publicly available replication package [^3] https://github.com/ICSE19-FAST-R/FAST-R provided by its authors to ensure a fair and consistent comparison.
4.5. Evaluation Metrics
Fault Detection Rate (FDR). We use the FDR to assess the effectiveness of RTM, which is defined as follows:
(10) $FDR=\frac{F^{{}^{\prime}}}{F}$
where $F$ is the total number of unique faults detected by the test suite before minimization, $F^{{}^{\prime}}$ is the total number of unique faults detected by the test suite after minimization. FDR ranges from 0 to 1, where 0 indicates that the minimized test suite detects no faults, while 1 signifies that it successfully detects all faults.
Requirement Coverage. Although RTM ensures full coverage of all requirements, one of the selected baseline approaches, i.e., FAST-R, can not guarantee complete coverage when ensuring the minimization budget. To provide a more comprehensive comparison, we also evaluate the requirement coverage of the baseline approaches. This is calculated as the ratio of the number of requirements covered by the minimized test suite to the total number of original requirements:
(11) $Coverage=\frac{|R^{\prime}|}{|R|}$
where $R^{\prime}$ represents the set of requirements covered by the minimized test suite $T^{\prime}$, and $R$ denotes the original set of requirements. The coverage value ranges from 0 to 1, where 0 indicates that the minimized test suite covers no requirements, while 1 signifies full coverage of all requirements.
4.6. Experiment Environment
All experiments were conducted on a computer provided by our industry partner, running Windows 10, equipped with an Intel Core i7-11850H CPU at 2.5 GHz and 32 GB of RAM. Due to data privacy considerations, we were restricted to performing our experiments solely on this laptop, which limited the choice of text representation techniques available to us.
5. Results
5.1. Performance of RTM under different configurations (RQ1)
Approach. To address RQ1, we systematically evaluated the performance of RTM under various configurations. Specifically, under a mid-range minimization budget ($50\%$), we ran RTM with three different preprocessing strategies (as described in Section 3.1), and applied both word-level and sentence-level embedding techniques to transform test cases into vector representations (Section 3.2). For measuring test case similarity, we compared several widely-used similarity metrics, including cosine similarity, Euclidean distance, and WMD. In addition, we investigated the impact of various GA initialization strategies on the overall performance of RTM by evaluating all three strategies introduced in Section 3.4.2.
Through a comprehensive grid search across these configurations, we identified the optimal configuration of RTM, i.e., the combination of preprocessing method, embedding technique, similarity metric, and initialization strategy that yields the best performance in terms of FDR.
Results. Table 1 presents the FDR of RTM across different configurations under 50% minimization budget. For each combination of initialization strategy and embedding technique, the best FDR is shown in bold, while the overall best result for each embedding technique is highlighted with a gray background.
Overall, we observe that sentence-level embeddings generally outperform word-level embeddings in terms of FDR. Among the sentence-level embedding techniques, TF-IDF combined with cosine similarity achieves the best overall performance, with a peak FDR of $86.09\%$ under $PM3$ and Init Strategy 2. This suggests that even simple, sparse representations like TF-IDF can be highly effective for TSM, especially when the test cases consist of structured test steps that mainly differ in variable names and parameter values. When combined with appropriate preprocessing methods and initialization strategies, TF-IDF effectively captures the discriminative textual features in such test cases by leveraging word frequency. Similarly, Amazon Titan Text Embedding V2, LongT5, and USE also achieve competitive results, with the best Titan configuration reaching $81.59\%$ ($PM3$, Init Strategy 2, Euclidean distance), LongT5 achieving $78.50\%$ ($PM3$, Init Strategy 3, Euclidean distance), and USE achieving $78.05\%$ ($PM1$, Init Strategy 3, Euclidean distance).
For word-level embeddings, Word2Vec slightly outperforms FastText and GloVe, but all three show lower FDR compared to sentence-level embedding techniques. Their performance peaks are in the range of $75\%-76\%$, with Word2Vec achieving the highest result of $76.27\%$ under $PM1$ and Init Strategy 2.
Among similarity measures, Euclidean distance outperforms cosine similarity for sentence-level embeddings most of the time, suggesting that the absolute distance measured by Euclidean-based spatial separation is more effective for capturing test case diversity in the embedding space.
In terms of preprocessing, we observe that, for word-level embedding, $PM1$ and $PM2$ often yield better FDR results than $PM3$, indicating that preprocessing is required for using word-level embedding techniques. Word-level embedding is more sensitive to individual characters such as casing and punctuation, and thus proper preprocessing helps normalize these variations, enabling more accurate similarity computations. On the other hand, $PM1$ and $PM3$ tend to lead to higher FDR than using $PM2$ for sentence-level embeddings, meaning that less or no preprocessing can help these models better learn context information from test cases, without losing the key details that distinguish them.
Lastly, Initialization Strategies 2 and 3 generally lead to higher FDR compared to Initialization Strategy 1, particularly when combined with stronger embeddings and similarity measures. This confirms the importance of starting the evolutionary search with a well-distributed and diverse initial subset.
Answering RQ1: TF-IDF with cosine similarity yields the highest overall FDR ($86.09\%$) under Preprocessing Method $PM3$ and Initialization Strategy 2 at the $50\%$ minimization budget.
Table 1. FDR (%) across different configurations (50% minimization budget)
Similarity Init Strategy 1 Init Strategy 2 Init Strategy 3
Measure $PM1$ $PM2$ $PM3$ $PM1$ $PM2$ $PM3$ $PM1$ $PM2$ $PM3$
TF-IDF Cosine 80.59 80.32 80.41 84.36 83.00 86.09 85.41 81.91 83.09
Euclidean 82.32 81.14 83.27 84.91 83.68 85.14 85.05 84.41 85.18
USE Cosine 70.36 72.68 70.41 77.05 76.27 77.41 77.59 77.05 77.59
Euclidean 71.68 73.91 72.68 77.68 76.50 77.05 78.05 75.82 77.50
LongT5 Cosine 72.82 72.68 72.86 76.14 77.09 76.14 77.23 76.45 76.18
Euclidean 74.05 74.64 74.91 77.77 78.23 78.32 78.36 77.82 78.50
Amazon Titan Text Embedding V2 Cosine 76.55 72.00 76.05 81.50 74.77 80.86 79.05 75.82 79.18
Euclidean 78.45 72.27 77.95 80.82 75.86 81.59 79.91 75.59 79.86
Word2Vec WMD 69.36 71.23 68.45 76.27 74.00 75.00 75.68 74.27 75.82
GloVe WMD 65.64 71.00 66.18 73.59 75.50 74.41 74.09 75.00 72.55
FastText WMD 66.55 71.41 67.00 74.82 73.82 74.32 74.09 75.68 73.32
5.2. Performance of RTM compared with baselines (RQ2)
Approach. To address RQ2, we compared the performance of RTM with the best configuration against baseline approaches in terms of FDR, varying the minimization budget from $10\%$ to $90\%$ in $10\%$ increments. According to RQ1, we selected RTM using $PM3$ as preprocessing strategy, TF-IDF as the embedding technique, Initialization Strategy 2 for GA initialization, and cosine similarity for the similarity measure as the best-performing configuration.
We evaluated RTM against Random Minimization, which also employs Initialization Strategy 2, and two versions of FAST-R (Cruciani et al., 2019), namely the fixed-budget version and the adequate version. For the fixed-budget version, we varied the minimization budget from $10\%$ to $90\%$, in increments of $10\%$. For the adequate scenario, we first executed FAST-R to determine its resulting minimization budget. This budget was then applied consistently to both RTM and the random minimization for experimental comparison. Notably, due to the redundancy in our dataset, as few as $7\%$ (i.e., $54$ test cases, one test case for each requirement) of the test cases are sufficient to achieve $100\%$ requirement coverage.
Moreover, for comparison purposes, we calculated the theoretical best possible FDR for the test suite under different minimization budgets using an ILP approach. Specifically, we employed ILP to find the subset that (1) satisfies the minimization budget, (2) achieves $100\%$ requirement coverage, and (3) maximizes the number of covered faults.
Results. Figure 3 depicts the FDR for RTM, Random Minimization and the fixed-budget version of FAST-R, as well as the theoretical best possible FDR across minimization budgets from $10\%$ to $90\%$.
Overall, RTM consistently outperforms all other techniques across minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.7\%$, whereas Random Minimization and FAST++ yield approximately $65.5\%$ and $56.5\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 00 or $1$, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
As the minimization budget increases, the FDR of all techniques gradually improves and approaches the theoretical upper bound of $1.0$. Among the FAST-R variants, FAST-CS and FAST-all demonstrate the more competitive performance, while FAST++ and FAST-pw exhibit weaker results at lower budgets, indicating limited effectiveness in resource-constrained settings. Interestingly, Random Minimization consistently achieves higher FDR than FAST-R. This may be due to the distance metrics employed by FAST-R, which may not provide effective guidance for reducing similar test cases, whereas Random Minimization inherently preserves a greater diversity of test cases, leading to better FDR.
[…]
Table 2. FDR (%) and Requirement Coverage (%) Comparison at Adequate Scenario for FAST-R
Approach 7% Budget 90% Budget
FDR Coverage FDR Coverage
RTM 26.32 100.00 98.59 100.00
FAST++ 27.27 100.00
FAST-CS 25.95 100.00
FAST-pw 21.18 89.44
FAST-all 92.59 99.63
RM with constraint 25.91 100.00 96.36 100.00
RM without constraint 19.86 39.26 95.41 98.52
RM: Random Minimization
In summary, compared with the baselines, RTM achieves the highest FDR across all minimization budgets, making it a more reliable choice for practical TSM with strong fault detection guarantees. In addition, RTM ensures full requirement coverage while enabling pre-defined minimization budgets, two features not simultaneously supported by FAST-R. These capabilities are particularly crucial in the context of requirement-based testing, where covering all requirements is often a strict necessity. Moreover, the ability to ensure fixed-budget minimization is essential in resource-constrained environments, where testing time or execution cost must be carefully managed.
Answering RQ2: RTM consistently outperforms all baselines in both FDR and requirement coverage under both the fixed-budget and adequate versions.
[…]
Figure 3. Comparison of FDR across minimization budgets for RTM, FAST-R, and Random Minimization, alongside the theoretical upper bound.
5.3. Impact of test suites redundancy (RQ3)
Approach. In RQ3, we investigate the impact of test suite redundancy levels on the performance (i.e.,FDR) of TSM techniques. Evaluating TSM techniques necessarily requires considering the test suite redundancy level, and we aim to shed light on this aspect. As described in Section 4.2, for each redundancy level (ranging from $4.5$ to $11.5$ in increments of $0.5$), we employ the ILP algorithm and GA to generate $10$ diverse test suites, each ensuring full coverage of both requirements and faults. We then apply RTM and the baseline methods to each of these test suites, and report the average FDR across the $10$ test suites for each redundancy level.
Results. Figure 4 illustrates the FDR performance of RTM against baseline methods (FAST++, FAST-CS, FAST-all, FAST-pw, and Random Minimization with and without full requirement coverage constraint) across varying redundancy levels (ranging from $4.5$ to $11.5$ in increments of $0.5$) under seven distinct minimization budgets (from $30\%$ to $90\%$ in increments of $10\%$). Moreover, for each test suite, we calculate the theoretical best FDR for comparison purpose. Note that we cannot set minimization budgets below $30\%$ as some test suites are unable to achieve full requirement coverage under those constraints.
We observe that RTM consistently achieves superior FDR compared to all baseline techniques across all redundancy levels and minimization budgets. Specifically, RTM more closely aligns with the theoretical optimum and consistently outperforms all other techniques. The difference in terms of FDR between RTM and baselines is particularly significant at mid-range minimization budgets ($30\%$$60\%$).
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward $1$. Random Minimization (with and without requirement coverage constraint) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, Random Minimization with the requirement constraint consistently outperformed the unconstrained variant in terms of FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage serves as an effective heuristic for enhancing fault detection under budgetary constraints.
Moreover, the analysis reveals that redundancy level significantly influences FDR at lower minimization budgets (budgets below $60\%$), highlighting the importance of considering redundancy under tight budget constraints. Notably, higher redundancy combined with low minimization budgets tends to result in larger gaps between the FDR of all TSM techniques and the theoretical maximum. This discrepancy might be explained as under a very low minimization budget, the search space is severely restricted, making it challenging to assemble a diverse subset that both covers all requirements and maximizes fault coverage. Future research should account for redundancy levels when evaluating TSM techniques.
[…]
Figure 4. Comparison of FDR Across Varying Redundancy Levels for RTM and Baseline Approaches Under Seven Minimization Budgets
Answering RQ3: The test suite redundancy level has a strong effect, in general, on the achieved FDR of all TSM techniques. Further, RTM consistently outperforms all baseline techniques in terms of FDR across all redundancy levels and minimization budgets.
6. Threats to Validity
In this study, we identified the following threats to validity:
External Validity. The generalizability of our findings may be limited due to the specific domain and nature of the test suites provided by our industry partner. The performance of our approach might differ when applied to other domains or software systems with different requirement structures, test suite characteristics, or fault profiles. To mitigate this threat, we selected test suites that vary in size and redundancy levels to evaluate the robustness of our approach under diverse configurations. Furthermore, our approach is designed to be domain-agnostic, relying solely on textual information from requirements and test cases, which facilitates potential adaptation to other domains. Last, our industry partner is following standard development practices in the automotive domain, imposed by standards, that are similar to other safety-critical domains where functional safety is important. Future work includes evaluating our approach on publicly available datasets from different domains to further assess the generalizability.
Internal Validity. To ensure a fair and consistent comparison, we adopt the original replication package of FAST-R provided by its authors. This minimizes potential implementation biases or configuration inconsistencies that could otherwise affect the validity of comparative results.
Another threat to external validity arises from the choice of embedding techniques that is limited by the data privacy constraints of our industry partner. This limitation may have affected the overall performance of our approach. Nevertheless, to mitigate this threat, we evaluated seven different embedding methods within the bounds of our available computational resources. This broad comparison helps reduce concerns regarding internal validity, despite the constraints on embedding selection.
Conclusion Threats. To reduce the impact of randomness and increase the reliability of our results, we run each experiment ten times with different random seeds and report the average performance. This mitigates the risk of drawing conclusions based on outlier results or fluctuations due to stochastic components in the minimization process.
7. Related Work
A number of approaches have been proposed to support TSM, which can be classified in to three categories (Khan et al., 2018; Cruciani et al., 2019; Pan et al., 2023): greedy (Noemmer and Haas, 2019; Miranda and Bertolino, 2017), clustering (Coviello et al., 2018; Liu et al., 2011; Viggiato et al., 2022; Cruciani et al., 2019), and search-based (Zhang et al., 2019; Hemmati et al., 2013; Pan et al., 2023, 2024; Marchetto et al., 2017), plus hybrid combinations (Xia et al., 2021; Anwar et al., 2019; Yoo and Harman, 2010) thereof.
Greedy-based approaches follow a heuristic, step-by-step selection strategy, where the test case that provides the most benefit toward the objective is chosen at each iteration, until the given constraints (e.g., full requirement coverage or code path coverage) are satisfied. Miranda et al. (Miranda and Bertolino, 2017) and Noemmer et al. (Noemmer and Haas, 2019) both employed greedy heuristics for TSM based on statement coverage. Their approaches iteratively selected test cases to maximize code coverage, achieving reasonable FDRs while significantly reducing the number of executed test cases. Greedy-based approaches are computationally efficient and easy to implement but may suffer from suboptimal global performance due to their locally optimal decisions
Clustering-based approaches group similar test cases based on syntactic or semantic similarities, such as those derived from natural language descriptions or execution behavior. The goal is to reduce redundancy by selecting representative test cases from each cluster. These approaches are particularly useful in settings where test cases are written in natural language and code coverage information is unavailable. Viggiato et al. (Viggiato et al., 2022) proposed a comprehensive approach that combines text embedding, similarity measurement, and clustering to identify and remove similar test cases written in natural language, where they split requirement test cases into test steps and then calculate the similarity between test steps. Cruciani et al. (Cruciani et al., 2019) proposed FAST-R, a black-box TSM approach that leverages the source code of test cases without requiring execution or coverage information. FAST-R transforms test code into vector representations using a term frequency model, followed by dimensionality reduction via random projection. Similarly, Coviello et al. (Coviello et al., 2018) proposed a clustering-based approach using Hierarchical Agglomerative Clustering to perform TSM.
Search-based approaches formulate the minimization problem as an optimization task and utilize metaheuristic algorithms (e.g., genetic algorithms, simulated annealing) to explore the space of possible test subsets. These methods are capable of balancing multiple conflicting objectives, such as minimizing the number of test cases while maximizing fault detection capability and maintaining requirement coverage. Hemmati et al. (Hemmati et al., 2013) applied a search algorithm that minimizes test suites based on model-derived test case similarity. Similarly, Zhang et al.(Zhang et al., 2019) proposed UncerTest, a model-based test case minimization framework that leverages multi-objective search algorithms, including NSGA-II, MOCell(Nebro et al., 2009), and SPEA2 (Zitzler et al., 2001). ATM (Pan et al., 2023) is a black-box TSM approach that combines test code similarity analysis with evolutionary search. It represents test code using Abstract Syntax Trees and employs Genetic Algorithm and Non-Dominated Sorting Genetic Algorithm II (Meyarivan et al., 2002) as the underlying evolutionary algorithms. Furthermore, LTM (Pan et al., 2024) leverages LLMs (CodeBERT, UniXcoder, and CodeLlama) to compute similarity between test cases and employs a Genetic Algorithm to search for an optimal subset of test cases under a fixed budget. MORE+ (Marchetto et al., 2017) is a multi-objective TSM approach that uses NSGA-II to optimize structural, functional, and cost-related objectives, aiming to enhance fault detection while reducing redundancy and execution time. Although typically more computationally intensive than greedy and clustering approaches, search-based techniques support multiple minimization objectives and often achieve better performance for TSM.
Hybrid approaches integrate complementary methods into a unified framework to balance multiple objectives, including reducing test suite size, preserving fault detection capability, and satisfying requirement or coverage constraints. Xia et al. (Xia et al., 2021) applies clustering-based test suite reduction combined with evolutionary multi-objective optimization. Anwar et al. (Anwar et al., 2019) combined genetic algorithm and particle swarm optimization algorithm to optimize regression test suites. Yoo et al. (Yoo and Harman, 2010) employs a hybrid multi-objective genetic algorithm that integrates the efficient approximation of the greedy approach with the capability of population-based genetic algorithms to generate higher-quality Pareto fronts.
While promising, existing techniques primarily focus on code-based solutions for minimization. In contrast, our work addresses minimization in the context of requirement-driven testing where test cases are specified in natural language and are thus introducing distinct challenges. Moreover, current approaches do not explicitly support TSM under a fixed budget while simultaneously enforcing full requirement coverage as a hard constraint. To the best of our knowledge, this is also the first work to investigate the impact of redundancy levels on fault detection capability under such constraints, providing valuable insights for future research and practical applications. Note that we retained FAST-R as one of the baselines though it cannot simultaneously satisfy both the fixed-size minimization budget and the full coverage adequacy constraint. This is because FAST-R can be applied in two scenarios (i.e., fixed-size budget and adequate), which enables experimental comparison with our approach.
8. Conclusion
In this paper, we presented RTM, a novel and practical approach for test suite minimization guided by requirements coverage. As it is often the case in domains where functional safety is crucial and must be demonstrated through adherence to standards, test cases are specified in natural language (test steps), are traced to requirements and are implemented manually.
By combining advanced text embedding techniques, multiple similarity metrics, and a Genetic Algorithm (GA) initialized using multiple strategies, RTM effectively reduces test suite size while strictly preserving full requirement coverage under a fixed budget constraint. Its evaluation on an industrial automotive dataset demonstrates its superior performance in Fault Detection Rate (FDR) across various minimization budgets. In comparison with multiple baselines, RTM consistently achieved significantly higher FDR, confirming its effectiveness in an industrial setting. Furthermore, our exploration of test suite redundancy levels revealed that higher redundancy significantly improves FDR outcomes, while higher redundancy combined with low minimization budgets tends to significantly increase the risk of undetected faults, thus highlighting the importance of considering redundancy characteristics in test suite minimization.
In conclusion, our results establish RTM as a scalable and reliable minimization solution for requirement-based testing (validation) in safety-critical systems. In the future, we plan to apply RTM to a broader range of application scenarios to further evaluate and enhance its generalizability.
Acknowledgements.
This work was supported by a research grant from Aptiv, as well as by the Discovery Grant and Canada Research Chair programs of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Research Ireland grant 13/RC/2094-2.